### v1

# CH2I2_UV_IR_CEI_SM

Abstract
Femtosecond structural dynamics of diiodomethane ($\mathrm{CH_{2}I_{2}}$) triggered by ultraviolet (UV) photoabsorption at 290 nm and 330 nm are studied using time-resolved coincident Coulomb explosion imaging driven by a near-infrared probe pulse. We map the dominant single-photon process, the cleavage of the carbon-iodine bond producing rotationally excited $\mathrm{CH_{2}I}$ radical, identify the contributions of the three-body ($\mathrm{CH_{2}+I+I}$) dissociation and molecular iodine formation channels, which are primarily driven by the absorption of more than one UV photon, and demonstrate the existence of a weak reaction pathway involving the formation of short-lived transient species resembling iso-$\mathrm{CH_{2}I-I}$ geometries with the slightly shorter I-I separation compared to the ground-state $\mathrm{CH_{2}I_{2}}$. These transient molecular configurations, which can be separated from the other channels by applying a set of conditions on the correlated momenta of three ionic fragments, are formed within approximately 100 fs after the initial photoexcitation and decay within the next 100 fs.
preprint: AIP/123-QED
I Introduction
The interaction between light and matter is a common and essential aspect of many fundamental processes in nature[1, 2], including photosynthesis[3, 4, 5], vision[6, 7, 8, 9], vitamin D synthesis[10, 11, 12], DNA repair[13, 14, 15], and various atmospheric reactions[16]. Among the latter, UV-induced photodissociation reactions of halogenated alkanes are significant sources of reactive halogens, which have a considerable impact on environmental and atmospheric chemistry. Iodine, one of the halogens, plays diverse roles in chemistry, serving as a fundamental element in human health and bio-chemistry, a useful catalyst in organic synthesis [17], and a major contributor to the destruction of ozone molecules [18]. Due to its strong absorption of sunlight in a broad range of UV wavelengths, diiodomethane ($\mathrm{CH_{2}I_{2}}$), a polyhalogenated alkane, is a major source of highly reactive iodine molecules influencing tropospheric chemistry and the marine boundary layer[19].
Numerous studies have been published on the UV-induced photochemistry of $\mathrm{CH_{2}I_{2}}$, which have reported the primary cleavage of one of the C-I bonds and the formation of $\mathrm{CH_{2}I}$ and $\mathrm{I(^{2}P_{3/2})}$ or $\mathrm{I^{*}(^{2}P_{1/2})}$ photoproducts [20, 21, 22]. It has been concluded that $\mathrm{I(^{2}P_{3/2})}$ and $\mathrm{I^{*}(^{2}P_{1/2})}$ are obtained via photodissociation from the $\mathrm{1B_{1}}$ and $\mathrm{2B_{1}}$ states, respectively. Due to possible couplings between $\mathrm{2B_{1}}$ and higher-lying electronic states, $\mathrm{I(^{2}P_{3/2})}$ could also be produced via non-adiabatic dissociation from the $\mathrm{2B_{1}}$ state [22, 23]. Baughcum et al. [22] reported that the ratio of the yield of $\mathrm{I^{*}(^{2}P_{1/2})}$ and $\mathrm{I(^{2}P_{3/2})}$ decreases as the excitation wavelength increases, with the quantum yield for $\mathrm{I^{*}(^{2}P_{1/2})}$ production dropping from 0.46 at 248 nm to 0.25 at 308 nm. Using ions generated by a (2+1) resonance enhanced multiphoton ionization process, Xu et al.[23] measured the translational kinetic energy distributions of both $\mathrm{I(^{2}P_{3/2})}$ and $\mathrm{I^{*}(^{2}P_{1/2})}$ fragments in the wavelength range of 277-305 nm and concluded that the $\mathrm{CH_{2}I}$ co-fragment is produced with significant internal excitation, with approximately 80% of the total available energy being partitioned into the internal energy of the $\mathrm{CH_{2}I}$ fragment.
Reid et al.[24, 25] later drew attention to the importance of isomerization of halogenated alkanes into iso-haloalkanes upon UV absorption as a major pathway leading to the production of molecular halogens. The photoisomerization of $\mathrm{CH_{2}I_{2}}$ to iso-$\mathrm{CH_{2}I_{2}}$ has been reported to be an extremely efficient process upon UV absorption in the solution phase and in cages[26, 27, 28]. Since it has been suggested that the dominant isomerization mechanisms in these studies are driven by the interaction with solvent molecules or cage, it is helpful to examine the process in isolated gas-phase molecules, free from such interactions. Borin et al.[29] reported observing photoisomerization of $\mathrm{CH_{2}I_{2}}$ in the gas phase upon 330 nm UV excitation by probing it with femtosecond transient absorption. The delayed rise in the transient absorption of $\mathrm{CH_{2}I_{2}}$ at 380 and 612 nm, approximately 35-90 fs after UV excitation, was attributed to the creation of a short-lived iso-$\mathrm{CH_{2}I_{2}}$. However, a recent ultrafast electron diffraction (UED) study, where Liu et al.[30] directly mapped the C-I cleavage upon 266 nm excitation, quantified the resulting increase of the C-I and I-I separations, and observed signatures of the rotation of the $\mathrm{CH_{2}I}$ radical after the C-I bond cleavage, did not mention any observation of iso-$\mathrm{CH_{2}I_{2}}$ formation.
Here, we employ Coulomb Explosion Imaging (CEI), a powerful method for determining the geometric structure of gas-phase molecules [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43], to investigate the structural dynamics of $\mathrm{CH_{2}I_{2}}$ upon UV photoabsorption. CEI has been shown to be a useful tool in identifying molecular isomers [35] and conformers [36] of polyhalogenated alkanes and other organic molecules [42], and even imaging the complete structure of halogenated alkanes [38, 39] and ring molecules [37, 40, 42]. When used as a time-resolved method in a pump-probe scheme, laser-induced CEI can measure wave-packets dynamics[44, 45, 46], visualize vibrational motions[47, 48, 46], and image photo-dissociation[49, 50, 51, 52] with high temporal resolution[53]. In particular, CEI provides multidimensional data crucial for identifying any transient structural changes in the molecule upon UV excitation, often with the help of simple classical modeling and computations.
Figure 1: Sketch depicting the pump-probe process and the different reaction pathways after photoexcitation of $\mathrm{CH_{2}I_{2}}$ (left) by the ultraviolet (UV) pump pulse. The products and possible intermediates (middle) are ionized to five-fold final charge state by the intense near-infrared (NIR) probe pulse, and the ionic fragments are detected in coincidence.
In this work, we investigate the dynamics of diiodomethane upon UV excitation using ionization and Coulomb explosion by an intense strong-field near-infrared (NIR) probe pulse. The primary objective of this time-resolved study is to explore possible signatures of the intramolecular photoisomerization process in $\mathrm{CH_{2}I_{2}}$, guided by the findings reported by Borin et al. [29]. We analyze the delay-dependent angular correlations between the momenta of the ions, the total kinetic energy release (KER), and the kinetic energies (KE) of individual ionic fragments to differentiate the possible photoisomerization channel from the direct two-body break-up channel that leads to $\mathrm{CH_{2}I+I}$ (or $\mathrm{I^{*}}$) products. We also identify the three-body ($\mathrm{CH_{2}+I+I}$) dissociation and the I2 formation channels, which primarily occur after the absorption of multiple UV photons. A schematic of the pump-probe experiment and the different reaction pathways after UV photoexcitation is shown in Fig. 1.
II Methods
II.1 Experimental
Figure 2: Schematic of the pump-probe setup and the coincident ion momentum imaging system. The diagram depicts the propagation of the ultraviolet (UV) pump and near-infrared (NIR) probe pulses into the experimental chamber. Both pulses are colinearly directed and focused onto a cold supersonic molecular jet containing diiodomethane. The setup is integrated with a double-sided velocity map imaging (VMI) spectrometer, operated in multi-ion coincidence mode. Ions are detected in coincidence using a time and position-sensitive detector comprising a microchannel plate (MCP) and delay line anode assembly. It should be noted that in this particular experiment, the electron detector is not utilized.
Figure 2 illustrates the experimental setup for the UV-pump and NIR-probe CEI experiment. The femtosecond laser system comprises of a 5-mJ, 3-kHz Coherent Legend Elite DUO Ti:Sapphire laser coupled to a Light Conversion TOPAS Prime optical parametric amplifier (OPA). The output of the Ti:Sapphire laser is split evenly into two independently compressed beam paths: one driving the OPA, and the other providing near-Fourier-transform-limited NIR pulses with a full width at half maximum intensity (FWHM) duration of 26 fs and a bandwidth of 60 nm at a central wavelength of 810 nm. In this experiment, the OPA was set to produce pulses at a central wavelength of 290 nm and 330 nm, with a 5-nm bandwidth, pulse duration of 70 fs (FWHM), and pulse energy up to 15 ${\mu}$J. The UV and NIR pulses were recombined using a 45-degree recombination mirror that transmitted the NIR pulses and reflected the UV pulses, and the combined beam was directed collinearly into the vacuum chamber through a 1 mm calcium fluoride $\mathrm{(CaF_{2})}$ window. The delay between NIR and UV pulses was adjusted via a computer-controlled optical delay stage in the NIR arm. Inside the vacuum chamber, the laser pulses are focused using a normal-incidence spherical mirror with 75-mm focal length and UV enhanced aluminium coating.
Diiodomethane, which is liquid at room temperature, is introduced into the vacuum chamber as a supersonic molecular beam expanded through a 30-${\mu}$m nozzle using helium at 3 psig as carrier gas, which is collimated by a skimmer with a 500-${\mu}$m diameter opening. A second differential pumping stage in the molecular beam, separated from the interaction chamber by a 700-${\mu}$m aperture, and a beam dump comprising of two differential pumping stages ensure that the base vacuum in the interaction chamber stays around $\mathrm{10^{-10}}$ mbar when the molecular beam is operating.
Ions produced by the interaction between the focused laser pulses and the molecular beam are collected using a double-sided velocity map imaging (VMI) spectrometer. This apparatus is similar to the one employed in the CAMP end-station for free-electron laser experiments [54, 55], while using position-sensitive delay-line detectors (Roentdek DLD80 and HEX80) for the coincident detection of multiple electrons and ions, as described for a similar setup by Ablikim et al. [56]. For the experiment described here, only the ion detector is used. The three-dimensional momentum vectors are obtained from the ions’ time of flight and hit positions as described by Lam et al.[42] and Ablikim et al. [56]. From these momentum vectors, we determine the KER and the angular correlations between the ionic momenta of all the coincidence channels that are of interest.
II.2 Coulomb explosion simulation
To guide the search for a possible isomer and identify the different reaction channels, we performed classical Coulomb explosion simulations for the molecular geometries corresponding to the $\mathrm{CH_{2}I_{2}}$ equilibrium geometry, the iso-$\mathrm{CH_{2}I_{2}}$, and the C-I bond cleavage pathway including rotation of the $\mathrm{CH_{2}I}$ intermediate. The simulations are based on solving Newton’s equations of motion for three-point charges located at the positions of the carbon and iodine atoms. To match the experimental observables, the $\mathrm{CH_{2}}$ fragment was kept intact. It is treated as a point particle in the simulations. Furthermore, the ionization and fragmentation process induced by the probe pulse was assumed to be instantaneous (i.e., instantaneously breaking both C-I bonds and resulting in a charge q=1 on the carbon atom and q=2 on each of the two iodine atoms), and the repulsion between the three fragments was treated as purely Coulombic. From prior experience, the calculations based on these assumptions typically overestimate the resulting KER but reproduce the experimentally observed momentum correlations [57, 58, 59, 60, 35, 36, 38, 42].
The optimized structures of the molecule in the equilibrium geometry of the neutral electronic ground state and the isomer geometry were taken from Borin et al. [29]. For each case, an ensemble of 5000 geometries was generated by randomly varying the initial atomic positions within a radius of 0.1 Å around the equilibrium geometries and the initial velocities by up to $\mathrm{2x10^{-4}}$ a.u., respectively. These parameters were chosen empirically in order to approximately match the width of the fragment kinetic energy and momentum distributions observed in the experiment. In the simulation for the dissociation channel, a rotational period of 300 fs (obtained experimentally) was used for the $\mathrm{CH_{2}I}$ intermediate, and the translational KEs used for the ion pairs were obtained from the experimental asymptotic KER for the respective channels.
III Results and discussions
While the main objective of this work is to investigate the transient photoisomerization of diiodomethane ($\mathrm{CH_{2}I}$-$\mathrm{I}$) after UV photoabsorption, the majority of the photoexcited molecules undergo direct dissociation following C-I bond cleavage, and the first task is therefore to identify the signatures of these competing reaction pathways. The intense NIR probe pulse ionizes both, the unpumped and the photoexcited molecules, resulting in multiple channels with different final charge states and different ionic fragments (see Figs. S1 - S4 of the Supplementary Material (SM) for ion time-of-flight (ToF) mass spectra and multi-ion coincidence spectra). The following discussion focuses on the analysis of the $\mathrm{CH_{2}^{+}+I^{2+}+I^{2+}}$ coincidence channel because, as discussed in detail in[61], this breakup channel contains minimal contribution from sequential fragmentation, which is prominent for channels with lower total charge states (see Fig. S5 and Fig. S6 of the SM), where it complicates the analysis. In Fig. 3a, the ion yield of this coincident channel is plotted as a function of the KER and the pump-probe delay.
Figure 3: (a) $\mathrm{CH_{2}^{+}+I^{2+}+I^{2+}}$ ion coincidence yield as a function of pump-probe delay and KER. Positive delays correspond to the NIR probe pulse arriving after the UV pump pulse. A step size of 15 fs and 100 fs was used for delays up to and beyond 800 fs, respectively. (b) KER spectra of the same coincidence channel recorded at a fixed delay of 13.5 ps. The data shown here were recorded at a pump wavelength of 290 nm. The equivalent plots for this and all the following figures with the data recorded at 330 nm are shown in Section II of the SM.
Figure 4: (a) $\mathrm{CH_{2}^{+}+I^{2+}+I^{2+}}$ ion coincidence yield as a function of KER and angle between the momentum vectors of the two iodine dications for a fixed pump-probe delay of 13.5 ps. Four distinct regions are marked by colored ovals corresponding to different contributions to the coincidence ion yield: Black: Coulomb explosion of bound molecules in or near the equilibrium geometry. Red: C-I cleavage and dissociation to $\mathrm{CH_{2}I}$ + $\mathrm{I}$. Green: Molecular iodine ($\mathrm{I_{2}}$) formation after UV absorption. Blue: UV-induced three-body dissociation into $\mathrm{CH_{2}}$ + I + I. (b) Coulomb explosion simulations of $\mathrm{CH_{2}I_{2}}$ for the same $\mathrm{CH_{2}^{+}+I^{2+}+I^{2+}}$ channel observed in the experiment, for different molecular geometries: the ground-state equilibrium geometry, the iso-$\mathrm{CH_{2}I_{2}}$ geometry according to Borin et al. [29], and for the three dissociation processes that are identified and marked in (a).
Three distinct features are visible in this plot: (i) a horizontal band in the KER range between 30 and 40 eV, which corresponds to the Coulomb explosion of bound (i.e. non-dissociating) molecules that may or may not have absorbed a UV photon and that were five-fold ionized by the probe pulse; (ii) a curved feature whose KER decreases with increasing pump-probe delay and which emerges from the horizontal band near time-zero, reaching a KER of approximately 15-18 eV at a delay of 2 ps and 9-14 eV at a delay of 13.5 ps (see Fig. 3b); and (iii) another curved feature whose KER decreases even faster than that of feature (ii), reaching a KER of approximately 8-10 eV at a delay of 2 ps and 1-5 eV at a delay of 13.5 ps, at which point the KER no longer changes with delay. Features (ii) and (iii) originate from the molecules that dissociated upon UV photoabsorption, with feature (ii) corresponding to two-body dissociations into $\mathrm{CH_{2}I}$ and $\mathrm{I}$ or $\mathrm{CH_{2}}$ and $\mathrm{I_{2}}$, while feature (iii) corresponds to direct three-body dissociation by the UV pulse into $\mathrm{CH_{2}}$, I and I. While these assignments can be made with good confidence based on the asymptotic KERs of both channels, further insights and confirmation are obtained when inspecting the angular correlation between the momentum vectors of the ionic fragments[61].
Figure 4a shows the coincidence ion yield at the asymptotic delay of 13.5 ps as a function of the KER and the angle between the momentum vectors of two iodine dications. The most intense feature at a KER of 30 - 40 eV (with a smaller side-peak between 20 - 30 eV) and relatively well defined ($\mathrm{I^{2+}}$, $\mathrm{I^{2+}}$) angle centered around 150 - marked by the black ellipse in Fig. 4a - stems from the Coulomb explosion of bound molecules in or near their equilibrium geometry, as confirmed by the data taken without the UV pulse present (see Fig. S6 in the SM). At a KER of approximately 15 eV (red ellipse), an angularly broad feature spanning 60-180 degrees in the ($\mathrm{I^{2+}}$, $\mathrm{I^{2+}}$) angle can be attributed to C-I cleavage and dissociation into $\mathrm{CH_{2}I}$ + $\mathrm{I}$ induced by a single-UV-photon absorption. The broad spread in angle results from high rotational excitation of the $\mathrm{CH_{2}I}$ fragment due to the torque imparted from the C-I bond cleavage, as also observed in other dihalomethanes [62, 63, 64, 65, 66, 67].
Two additional features, each of them with approximately half the number of events as in the red ellipse discussed above, are prominent in Fig. 4. First, a rather localized spot peaked at a KER slightly below 20 eV and an ($\mathrm{I^{2+}}$, $\mathrm{I^{2+}}$) angle close to 180° (green ellipse) can be uniquely attributed to $\mathrm{I_{2}}$ formation after UV excitation (i.e., dissociation into $\mathrm{CH_{2}}$ + $\mathrm{I_{2}}$) since both the KER and the back-to-back emission of the two iodine ions is consistent with a Coulomb explosion of $\mathrm{I_{2}}$. Second, a clearly separated contribution at low KER (blue ellipse) is attributed to the UV-induced three-body dissociation into $\mathrm{CH_{2}}$ + I + I. Both of these features result mainly from multi-photon excitation by the UV pulse. This assignment is based on (i) the energy required to trigger three-body dissociation (4.8 - 5 eV [68, 69]), (ii) earlier experimental work that reported rather low quantum yield of $\mathrm{I_{2}}$ elimination upon single-photon excitation [68], and (iii) on the observed dependence on the UV power. A detailed analysis of the UV power dependence of the different channels for pump wavelengths of 290 nm and 330 nm is presented in Fig. S12 of the SM.
The results of the classical Coulomb explosion simulations for the neutral ground-state geometry and the iso-$\mathrm{CH_{2}I_{2}}$ equilibrium geometry are shown in Fig. 4b, along with the simulation results for the $\mathrm{CH_{2}I-I}$, $\mathrm{CH_{2}-I_{2}}$ and $\mathrm{CH_{2}-I-I}$ dissociation channels at 13.5 ps pump-probe delay. Apart from overestimating the measured KER, most likely due to the assumption of a purely Coulombic potential, the simulated results qualitatively match the experimental observations for all channels, validating the use of these simulations in identifying the observables corresponding to the possible formation of the isomer.
Although the simulations show a subtle difference in the KER of the equilibrium geometry and the isomer, this distinction can be expected to be much less pronounced in the experiment, where a significant spread in the KER is observed even when no UV pulse is present. However, the simulations also predict that the Coulomb explosion of the isomer will lead to some ion yield at larger angles between the ($\mathrm{I^{2+}}$, $\mathrm{I^{2+}}$) momenta than those realized by the Coulomb explosion of the equilibrium geometry. In the experiment, essentially no events are observed with ($\mathrm{I^{2+}}$, $\mathrm{I^{2+}}$) angles larger than 160° when no UV pulse is present (see Fig. S6 in the SM), but a transient signal with high KER is observed in this region for pump probe delays between 0 and 250 fs (see Supplementary Movie 1). By selecting only those events in the experimental data where the ($\mathrm{I^{2+}}$, $\mathrm{I^{2+}}$) angle is larger than 160°, we can thus almost completely suppress the contribution from molecules in the equilibrium geometry.
Figure 5: Coulomb explosion simulations of $\mathrm{CH_{2}I_{2}}$ for the $\mathrm{CH_{2}^{+}+I^{2+}+I^{2+}}$ breakup channel for different molecular geometries: the ground-state equilibrium geometry, the iso-$\mathrm{CH_{2}I_{2}}$ geometry according to Borin et al. [29], and for a specific geometry during the C-I dissociation process when the $\mathrm{CH_{2}}$-I-I angle in the dissociating molecule is the same as in iso-$\mathrm{CH_{2}I_{2}}$ due to the rotation of the $\mathrm{CH_{2}I}$ radical. The two panels show the distribution of these events (a) as a function of the sum KE of the two iodine fragments and angle between the two $\mathrm{I^{2+}}$ momentum vectors, and (b) normalized yield of the KE sum of the two $\mathrm{I^{2+}}$ ions integrated for all angles.
However, from the experimental data, it is evident that the three dissociation pathways - (i) $\mathrm{CH_{2}I}$-I two-body dissociation with $\mathrm{CH_{2}I}$ rotation, (ii) $\mathrm{CH_{2}}$-I-I three-body dissociation and (iii) molecular $\mathrm{I_{2}}$ formation - could also contribute to the region with the ($\mathrm{I^{2+}}$, $\mathrm{I^{2+}}$) angle larger than 160° (at least at large delays), as seen in Fig. 4a. Contributions from these pathways need to be filtered out to identify possible signatures of isomerization to $\mathrm{CH_{2}I}$-I. The simulations in Fig. 5 show that after Coulomb explosion, the sum of the KEs of the two $\mathrm{I^{2+}}$ fragments is larger for the iso-$\mathrm{CH_{2}I_{2}}$ molecules than that of the $\mathrm{CH_{2}I_{2}}$ molecules that undergo direct dissociation into $\mathrm{CH_{2}I}$ and I via C-I cleavage. Furthermore, in the cases of three-body dissociation or molecular iodine formation, the KE of the methyl ion decreases rapidly with pump-probe delay as the $\mathrm{CH_{2}}$ fragment quickly moves away from the two iodine atoms or the $\mathrm{I_{2}}$ molecule. Therefore, the sum of the KEs of the two $\mathrm{I^{2+}}$ fragments as well as the KE of the $\mathrm{CH_{2}}$ fragment can serve as additional parameters to discriminate the different channels.
Since it is very likely that any isomer-like geometries, were they to be formed as predicted by Borin et al., would be visited quickly due to the large amount of internal energy deposited into the molecule by the UV photon, we will concentrate our search on the smaller delays below 1 ps, where the isomer-geometry should appear as an additional contribution in the region corresponding to the bound molecules. Note that we have, nonetheless, also searched at larger pump-probe delays but have not found any statistically significant indications of isomer-like geometries at larger delays.
Figure 6: Kinetic energy of the $\mathrm{CH_{2}^{+}}$ fragment as a function of the pump-probe delay, with only those coincidence events shown where the angle between the two $\mathrm{I^{2+}}$ momentum vectors is greater than 160. The region below the blue line, at 13 eV kinetic energy of methyl cation, corresponds to events from $\mathrm{CH_{2}}$-I-I three-body dissociation and molecular $\mathrm{I_{2}}$ formation pathways (see text). The region of interest, above 13 eV, is where any events from photoisomerization following UV excitation would appear, along with the contribution from two-body dissociation producing I (or $\mathrm{I^{*}}$) and rotationally excited $\mathrm{CH_{2}I}$.
In Fig. 6, the KE distribution of the methyl cation ($\mathrm{CH_{2}^{+}}$) is shown as a function of the pump-probe delay up to 800 fs and only for those events where the ($\mathrm{I^{2+}}$, $\mathrm{I^{2+}}$) angle is larger than 160°. Based on the reasoning outlined in the previous paragraph, we can conclude that the events with quickly decreasing KE, below the blue line in Fig. 6, correspond to either $\mathrm{CH_{2}}$-I-I three-body dissociation or molecular iodine formation. The events above the blue line at 13 eV mainly correspond to dissociation of $\mathrm{CH_{2}I_{2}}$ into $\mathrm{CH_{2}I}$ and $\mathrm{I}$ (with subsequent rotation of the $\mathrm{CH_{2}I}$ fragment) as well as any possible formation of iso-$\mathrm{CH_{2}I_{2}}$, with only spurious contributions from the three-body dissociation and molecular iodine formation pathways at small delays below 200 fs.
Next, we select only those events with kinetic energy of $\mathrm{CH_{2}^{+}}$ ionic fragment above 13 eV from Fig. 6 and plot the sum of the KEs of the two iodine fragments as a function of pump-probe delay in Fig. 7(a). The plot reveals two contributions: an intense feature with a sum KE between 10 and 17 eV (marked by the red rectangle), and a much weaker feature at a higher sum KE, marked by the green rectangle, which contains approximately 7% of the events in the intense feature. The projection in panel (b) reveals that these two contributions (blue line) are shifted toward lower and higher energies, respectively, compared to the sum KE of the unpumped molecules (red dotted line). The integrated yield inside the green region of interest as a function of the pump-probe delay is shown as green line in Fig. 7(c). Based on the simulations shown in Fig. 5(b), we assign the events with the higher sum KE to molecular geometries resembling the iso-$\mathrm{CH_{2}I_{2}}$ structure predicted by Borin et al., which have a slightly smaller I-I distance than geometries resulting from the dissociation of $\mathrm{CH_{2}I_{2}}$. Fig. 7(c) shows that the occurrence of these structures peaks at a delay of approximately 120 fs and quickly decays again on a similar time scale. Analysis of the pump-probe data recorded at a pump wavelength of 330 nm, following the exact same steps as described above (see section II in SM), yields the blue dashed line in Fig. 7(c), which shows the same behavior as the data at 290 nm except for possibly a subtle shift towards slightly longer delays, which is at the borderline of statistical significance for the present data set.
Figure 7: (a) Coincidence ion yield as a function of pump-probe delay and KE sum of the two $\mathrm{I^{2+}}$ fragments, for only those events where the angle between the two $\mathrm{I^{2+}}$ momentum vectors is greater than 160° and the KE of the $\mathrm{CH_{2}^{+}}$ fragment is greater than 13 eV (above the blue line in Fig. 6). The two regions of interest are described in the text. (b) Projection of the coincident ion yield in panel (a) on the KE axis. The corresponding KE sum distribution for "unpumped" molecules (NIR probe pulse arrives before the UV pump pulse) is shown as a red dotted line. (c) Projection of the coincident ion yield in the green ROI in (a) on the pump-probe delay axis (green), compared to the similarly analyzed yield at a pump wavelength of 330 nm (blue).
Figure 8: Energy sharing between the two $\mathrm{I^{2+}}$ fragments, integrated over the pump-probe delays shown in Fig. 7, for only those events where the angle between the two $\mathrm{I^{2+}}$ momentum vectors is greater than 160° and (a) the KE of the $\mathrm{CH_{2}^{+}}$ fragment is greater than 13 eV (above the blue line in Fig. 6); or (b) the KE of the $\mathrm{CH_{2}^{+}}$ fragment is less than 13 eV (below the blue line in Fig. 6). The red and green rectangles in panel (a) correspond to events originating from the respective rectangular regions in Fig. 7.
Additional information on the geometry of the transient species can be obtained by considering the energy sharing between the two detected $\mathrm{I^{2+}}$ fragments, as shown in Fig. 8. Symmetric geometries, such as the original $\mathrm{CH_{2}I_{2}}$ geometry, or any molecules excited to symmetric vibrational modes, would lead to equal energy sharing, whereas strongly asymmetric geometries such as the predicted iso-$\mathrm{CH_{2}I_{2}}$ species, would lead to a pronounced asymmetry in the energy sharing. From Fig. 8(a), it is evident that all the events surviving both the filters - large $\mathrm{I^{2+}-I^{2+}}$ angle and higher $\mathrm{CH_{2}^{+}}$ energy, have asymmetric energy sharing between the $\mathrm{I^{2+}}$ fragments, excluding, for example, the possibility that any of these events stem from a symmetric $\mathrm{I_{2}}$ formation channel. In contrast, symmetric geometries contribute strongly to the coincidence yield with low $\mathrm{CH_{2}^{+}}$ energy, as shown in Fig. 8(b), corroborating our previous conclusions that many of those events are due to $\mathrm{I_{2}}$ formation.
At this point, we would like to emphasize that although detailed analysis of the CEI data clearly shows the transient formation of $\mathrm{CH_{2}I-I}$ isomer-like geometries that are consistent with previous literature [29], the CEI experiment alone is not able to confirm this as an "isomerization" process, nor can the CEI experiment identify the electronic state of the molecule when it assumes these isomer-like geometries.
Whether or not the process observed and reported here can or should be referred to as "isomerization" cannot be answered by the experimental data alone and also depends on the exact definition of "isomerization" one chooses to use. Given the short time that the molecules spend in the isomer-like geometry, which is less than the timescale of the rotation of the $\mathrm{CH_{2}I}$ fragment that is produced by direct C-I bond cleavage, the process might be better described as a transient passage through isomer-like geometries rather than an actual trapping in the potential well of the isomer, but these terminologies and interpretations strongly rely on careful theoretical modeling, which is beyond the scope of the present paper.
IV Conclusion
We have applied time-resolved three-body Coulomb explosion imaging to study the UV-induced (290-nm and 330-nm) photochemical dynamics in gas-phase $\mathrm{CH_{2}I_{2}}$. In addition to the primary dissociation into $\mathrm{CH_{2}I+I}$ via direct cleavage of the C-I bond, the formation of molecular $\mathrm{I_{2}}$, and the three-body dissociation into $\mathrm{CH_{2}+I+I}$ (where the latter two processes are attributed to the absorption of more than one UV photon), we also observe the transient molecular configurations resembling iso-$\mathrm{CH_{2}I_{2}}$ geometries, which are formed within approximately 100 fs after the photoexcitation and completely decay within an additional 100 fs. Within the uncertainties of the experimental data, the relative yield, formation and decay time of the transient iso-$\mathrm{CH_{2}I_{2}}$ products appear to be similar at both excitation wavelengths employed in this study, suggesting that their formation does not depend strongly on the excitation wavelength.
Acknowledgements.
We gratefully acknowledge the technical staff of the J.R. Macdonald Laboratory for their excellent support of the experiments; former PhD students Dr. Seyyed Javad Robatjazi and Dr. Shashank Pathak for commissioning the double-sided VMI apparatus; and Prof. Brett Esry for insightful discussions throughout the project. This work was supported by the Chemical Sciences, Geosciences, and Biosciences Division, Office of Basic Energy Sciences, Office of Science, US Department of Energy, grants no. DE-FG02-86ER13491 and DE-SC0020276 (S.B.), and by the National Science Foundation grant no. PHYS-2409365 (A.S.V.).
Conflict of Interest
The authors have no conflicts to disclose.